OpenStudy (anonymous):

What are the explicit equation and domain for a geometric sequence with a first term of 4 and a second term of -8? (5 points) an = 4(-2)n - 1; all integers where n ≥ 0 an = 4(-2)n - 1; all integers where n ≥ 1 an = 4(-12)n - 1; all integers where n ≥ 1 an = 4(-12)n - 1; all integers where n ≥ 0

OpenStudy (anonymous):

PLEASE SOME ONE HELP

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi!

OpenStudy (anonymous):

do you mind helping me with this

OpenStudy (anonymous):

i believe that the answer is the third one but im not sure

OpenStudy (misty1212):

\[a,ar,ar^2,ar^3...\] \[4,4\times r=-8\]

OpenStudy (misty1212):

if \[4\times r=-8\] then \[r=\frac{-8}{4}=-2\]

OpenStudy (anonymous):

okay

OpenStudy (misty1212):

therefore \[a_n=4\times (-2)^{n-1}\]

OpenStudy (anonymous):

so is it a or b???

OpenStudy (misty1212):

it would be B

OpenStudy (anonymous):

okay thank you

OpenStudy (misty1212):

because \(n\geq 1\) is what they want as the domain

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]